If $5530 is invested in an account at 5% interest compounded monthly, how long will it take the investment to double? A=P (1+ r/m)^mt
Answer by Boreal(15235)

A=P (1+ r/m)^mt
A=2P, because it is doubling.
2P=P (1+ r/m)^mt
2=(1+ r/m)^mt
2=(1+.05/12)^12t
ln of both sides
ln2=12t ln (1.004166). Without rounding, divide ln 2 by the ln on the right and divided that answer by 12, and t will be the result.
t=13.89 years
using the rule of 70, 70/5, the interest rate, equals 14 years. The answer makes sense.